Strange attractors in periodically forced limit cycles

Lai-Sang Young Courant Institute, USA

We discuss the emergence of chaotic behavior in the form of horseshoes and strange attractors with SRB measures when certain simple dynamical systems are kicked at periodic time intervals. The settings considered include limit cycles and stationary points undergoing Hopf bifurcations.